Interactive LearningWare 18.1 offers some perspective on this problem. Two tiny spheres have the same mass and carry charges of the same magnitude. The mass of each sphere is 3.20 × 10-6 kg. The gravitational force that each sphere exerts on the other is balanced by the electric force. Determine the charge magnitude.
Answer #1

using equilibrium of forces

electric force = gravitational force

k q^2 / r^2 = G m^2 / r^2

9*10^9* q^2 = 6.67*10^-11* (3.2*10^-6)^2

q = 2.755*10^-16 C

=======

Comment in case any doubt will reply for sure . Goodluck